OpenStudy (anonymous):

A ferry makes an 80km trip down the river in 5 hours. The trip back against the current takes 8 hours. Fin the speed of the ferry and the current.

OpenStudy (ddcamp):

Let's create some variables to help: f - speed of the ferry c - speed of the current So, what's the total (or net) velocity when the ferry is moving with the current. Against the current?

OpenStudy (anonymous):

f+c and f-c

OpenStudy (ddcamp):

OK. We know that, both ways, it traveled 80km. One way, it took 5 hours, the other way took 8. Do you know the equation for average speed?

OpenStudy (anonymous):

D=RxT? Not sure..

OpenStudy (ddcamp):

Yeah, that's the one. Let's rearrange it a bit though: \[speed = \frac{distance}{time}\]

OpenStudy (ddcamp):

Using that, can you find the average velocity for each direction?

OpenStudy (ddcamp):

*average speed

OpenStudy (anonymous):

80/5 and 80/8

OpenStudy (ddcamp):

Yup! Those simplify to 16 and 10. So, when the ferry is traveling with the current, the speed is 16 km/h, and when it's traveling against the current, the speed is 10 km/h. Using these numbers, and the equations you found earlier, can you set up a system of equations?

OpenStudy (anonymous):

80=16f+c and 80=10f-c?

OpenStudy (anonymous):

oh wait nevermind

OpenStudy (anonymous):

Im not sure haha

OpenStudy (ddcamp):

Remember, the first equation (f+c) is the total speed when the ferry is travelling with the current. f-c is when it's travelling against. This leaves us with: f + c = 16 f - c = 10

OpenStudy (anonymous):

ohh alright, then we use elimination?

OpenStudy (ddcamp):

Whichever method is easiest for you.

OpenStudy (anonymous):

Ok awesome, thanks so much :) you were a big help